# StructBench: An Autogenerated Benchmark for Evaluating Large Language Model’s Ability in Structure-Rich Text Understanding

### v1

@@ Abstract
Given the substantial volumes of structured data held by many companies, enabling Large Language Models (LLMs) to directly understand structured text in non-structured forms could significantly enhance their capabilities across various business scenarios. To this end, we propose evaluation data generation method for assessing LLM’s ability in understanding the structure-rich text, which generates structured data of controllable complexity based on manually crafted question templates and generation rules. Building on this generation method, we introduce StructBench, a benchmark comprising 6,032 questions across 8 different structured languages and 29 specific tasks. Furthermore, considering human proficiency in rule-based tasks, we also present StructBench-Hard, which includes 3,016 questions designed to further examine the gap between LLMs and human performance. Results indicate that the best-performing LLM currently achieve an accuracy of 65.0% on StructBench-Hard, while human accuracy reaches up to 95.7%. Moreover, while fine-tuning using StructBench can enhance existing LLMs’ understanding of all structured languages, it does not necessarily improve performance across all task types. The benchmark and generation codes are open sourced in https://github.com/MikeGu721/StructBench
@@ 1 Introduction
In response, we propose a method for automatically generating evaluation data to assess models’ capabilities in understanding structure-rich text. This method is applied to 8 structured languages, as shown in Fig. 1, across 29 specific tasks, enabling data generation with controllable difficulty by adjusting the depth of structured nesting. Based on this method, we further introduce the Structure-Rich Text Benchmark (StructBench), a comprehensive benchmark with 6,032 samples designed to evaluate the proficiency of LLMs in deciphering embedded structures within input text. StructBench aims to determine whether LLMs understand raw structural tags, execute logical inferences based on the decoded semantics of these symbols, and construct their responses according to instruction requirements. Considering that humans excel at understanding structured expressions, we further propose StructBench-Hard, a subset of StructBench consisting of 3,012 samples with the longest questions, highlighting the disparity between LLMs and human understanding of structured data.
The experiments show that StructBench significantly challenges current LLMs. The highest accuracy achieved by LLMs using various prompt methods was 78.4%, dropping to 65.0% on StructBench-Hard. In contrast, human participants reached 95.7% accuracy on StructBench-Hard, underscoring the inadequacy of existing LLMs in processing complex structural information. Utilizing StructBench data enhances the capabilities of current closed-source LLMs and improves open-source LLMs by fine-tuning them with this benchmark. This finding confirms StructBench’s critical role in evaluating and optimizing model performance. However, more than using the training data from StructBench alone is required to surpass GPT4-Turbo’s performance on this test suite. Additionally, not all tasks can be improved solely by training on the relevant dataset. Therefore, StructBench provides a comprehensive evaluation framework for assessing LLMs’ ability to understand structured text, leaving room for further improvement.
We present the following contributions to our research: (1) We developed a method for generating evaluation data for eight structured languages. This method allows control over the complexity of the generated evaluation data by manipulating the quantity of structural nesting. (2) We introduce StructBench, a comprehensive benchmark comprising 6,032 samples across eight structured languages and 29 evaluation tasks. This benchmark is designed to assess LLMs’ capabilities in understanding structured data comprehensively. (3) We have conducted extensive experiments, including evaluations of existing large models’ capabilities in understanding structure-rich text and enhancements to the capacity of open-source LLMs in this area. Our findings indicate a significant gap between current LLMs and humans in understanding structure-rich text.
@@ 2.2 Structural Text Understanding Evaluation Datasets
The similar datasets dedicating to structure information understanding are scarce. GraphQA [9] is a benchmark for assessing LLMs on reasoning over graph data, composed of 7 simple graph tasks, with exploration on methods in encoding graph-structured data as text. Sui et al. [21] consists of 8 tasks on table-formatted input texts, i.e., input data organized in row-column structures, with studying in input design and prompt enhancement. Pasupat and Liang [17] exposed a dataset of 22,033 complex questions on Wikipedia tables and proposed a parsing algorithm for answering complex questions on semi-structured tables. For tabular data, Shwartz-Ziv and Armon [20] showed that deep models were not a good solution; Gorishniy et al. [12] designed two specialized baseline architecture for tabular data and concluded that universally superior solution does not exist. Wang et al. [27] is another graph-based problem solving benchmark, with graph structure described with natural language. Fatemi et al. [9] and Sui et al. [21] include comparison of encoding methods for structure information, i.e., graphs and tables, respectively, showing that LLMs performance varies on different encodings and the relatively vacancy on LLM’s understanding of structure-rich texts input.
Our dataset diverges from previous work in two significant ways. Firstly, our tasks exclusively concentrate on analyzing structural information and executing dependent inference, unlike prior approaches [16, 31] that employ structural data primarily as an auxiliary tool for classification or semantic comprehension. StructBench focus on symbolic structure by erasing content semantics from input texts, preserving only structural details, which renders the texts semantically obscure and the tasks semantically agnostic. Secondly, our benchmark extends beyond traditional graph reasoning and table information retrieval to include a broader spectrum of structure-rich text types. It also incorporates comprehensive structure-based inference tasks that are tailored to specific input formats.
3 StructBench Construction
@@ 3.3 Statistic Information
StructBench has assembled a dataset comprising 6,032 samples, with each of the 29 specific tasks for eight languages containing 208 samples. These tasks are divided into “simple” and “hard” categories based on data length and complexity, as depicted in Fig. 3(a). The dataset is partitioned into training and testing sets at a 9:2 ratio, which supports extensive training while also enabling the evaluation of model generalization on significant volumes of unseen data. Detailed statistics regarding the number of samples, lengths, and complexity levels across all tasks, languages and difficulties are detailed in Tab. 3 in Sec. C in Appendix.
@@ 4 Experiment Setup
To evaluate LLMs’ current capability of processing structure-rich text and executing dependent inference, we conducted a series of experiments using StructBench in various settings. Our study utilizes both prompt-based and finetuning methods to analyze the performance variations.
@@ 4.3 Finetune-based Method
We further explored the efficacy of finetuning OpenLLaMA 7B v1 using various subsets of data. Our finetuning configurations are described as follows: Base: Utilizes the core StructBench dataset, referred to as the base StructBench setting. 2-Shot: Implements a few-shot (2-shot) learning strategy within the StructBench framework. Rule Hint: Adapts StructBench by appending a new task-specific rule hint to each sample, effectively redefining the input text as a combination of the question, answer, and original input, with the answer delineated as a step-by-step procedure. Mix: is a variant of Base with half of samples replaced by samples from ShareGPT111https://huggingface.co/datasets/shibing624/sharegpt_gpt4.
@@ 5.1 Close Sourced LLMs’ Performance
In this section, we analyze the performance of closed-source LLMs in various tasks involving structure-rich text and multiple languages. Table 5 displays the overall performance on the complete StructBench dataset, while Table 2 compares both LLM and human performance specifically on the StructBench-Hard subset.
@@ 5.3 Case Study
We illustrate the assessment setup and specific evaluation methods of StructBench through two exemplary cases in Fig. 8.
@@ 6 Conclusion
We believe that the capability to directly interpret structural-rich text in a free-text format is an essential skill all LLMs require. In response, we have developed StructBench to evaluate this capability of LLMs. Our findings indicate that the proficiency of current LLMs in training on these structural-rich texts varies depending on user frequency, leading to markedly different outcomes when the same tasks are performed in various languages. Currently, LLMs’ understanding of structural-rich texts remains superficially tied to the training data, and these models lack a profound understanding of the structure itself. This deficiency becomes evident when LLMs encounter complex structures composed of common languages or need to parse structural-rich text by custom languages, resulting in significant performance degradation. Direct training on the dataset provided by StructBench effectively enhances the LLM’s understanding of structural-rich text. However, training solely on the StructBench dataset does not allow achieving GPT4-Turbo’s comprehensive capabilities on the test set, necessitating the collection of more extensive data and the optimization of training methodologies.
@@ 7 Limitation
This paper focuses on evaluating and enhancing the comprehension LLM’s abilities on structure-rich text by designing a benchmark called StructBench. However, StructBench includes only eight types of structured languages and encompasses a total of 29 different tasks. Given the vast array of actual structured languages and the myriad methodologies employed beyond these 29 types, StructBench can only partially represent the LLMs’ capacity to understand structure-rich text. Additionally, due to regional restrictions, we are unable to utilize some highly effective baseline LLMs, such as Gemini and Claude. Therefore, the conclusions drawn in this paper are based on the assumption that GPT-4 and GPT-4 Turbo represent the top-tier LLMs now.
@@ 8 Ethical Concern
Nature of StructBench Content: StructBench is primarily composed of structured language syntax and some nonsensical strings, which do not present potential ethical issues such as gender bias or racial discrimination.
@@
Objective Presentation of Experimental Results: The experimental results pertaining to StructBench objectively demonstrate the comprehension abilities of various large models on structure-rich text included in the benchmark. We have thoroughly validated the outputs and assessment details of the models to ensure that the entire evaluation adheres to the experimental setup and maintains objectivity.
@@
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Sec. 1
Did you describe the limitations of your work? [Yes] See Sec. 7
Did you discuss any potential negative societal impacts of your work? [Yes] See Sec. 8
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] See Sec. 8
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Sec. 3
- (b)
Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] See Sec. C
Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Sec. C
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
If your work uses existing assets, did you cite the creators? [Yes] See Sec. 4
Did you mention the license of the assets? [Yes] See Sec. 4
Did you include any new assets either in the supplemental material or as a URL? [Yes] See https://github.com/YHN-ice/StructBench
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See Sec. 3
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See Sec. 8
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Sec. A
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [Yes] See Sec. A
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] See Sec. A
@@ Appendix A Detail about Manual Works
This paper involves the manual works in writing Question Templates and the acquisition of human performance on StructBench-Hard. All manual works are carried out by the authors of this paper, so there is no payment for the works. The three authors collectively completed the writing and verification of 29 Question Templates, and all templates along with the dataset have been made publicly available on the same website. Moreover, each of the three authors provided responses to 500 identical questions in StructBench-Hard, with each author dedicating approximately 17 hours. Thus, the human performance results presented in Table 2 are calculated based on the average scores across these 1500 responses.
Appendix B Statistics to StructBench
The detailed statistic information about StructBench is illustrated in Tab. 3.

### v2

# StrucText-Eval: An Autogenerated Benchmark for Evaluating Large Language Model’s Ability in Structure-Rich Text Understanding

@@ Abstract
Given the substantial volumes of structured data held by many companies, enabling Large Language Models (LLMs) to directly understand structured text in non-structured forms could significantly enhance their capabilities across various business scenarios. To this end, we propose evaluation data generation method for assessing LLM’s ability in understanding the structure-rich text, which generates structured data of controllable complexity based on manually crafted question templates and generation rules. Building on this generation method, we introduce StrucText-Eval, a benchmark comprising 6,032 questions across 8 different structured languages and 29 specific tasks. Furthermore, considering human proficiency in rule-based tasks, we also present StrucText-Eval-Hard, which includes 3,016 questions designed to further examine the gap between LLMs and human performance. Results indicate that the best-performing LLM currently achieve an accuracy of 65.0% on StrucText-Eval-Hard, while human accuracy reaches up to 95.7%. Moreover, while fine-tuning using StrucText-Eval can enhance existing LLMs’ understanding of all structured languages, it does not necessarily improve performance across all task types. The benchmark and generation codes are open sourced in https://github.com/MikeGu721/StrucText-Eval
@@ 1 Introduction
In response, we propose a method for automatically generating evaluation data to assess models’ capabilities in understanding structure-rich text. This method is applied to 8 structured languages, as shown in Fig. 1, across 29 specific tasks, enabling data generation with controllable difficulty by adjusting the depth of structured nesting. Based on this method, we further introduce the Structure-Rich Text Evaluation Benchmark (StrucText-Eval), a comprehensive benchmark with 6,032 samples designed to evaluate the proficiency of LLMs in deciphering embedded structures within input text. StrucText-Eval aims to determine whether LLMs understand raw structural tags, execute logical inferences based on the decoded semantics of these symbols, and construct their responses according to instruction requirements. Considering that humans excel at understanding structured expressions, we further propose StrucText-Eval-Hard, a subset of StrucText-Eval consisting of 3,012 samples with the longest questions, highlighting the disparity between LLMs and human understanding of structured data.
The experiments show that StrucText-Eval significantly challenges current LLMs. The highest accuracy achieved by LLMs using various prompt methods was 78.4%, dropping to 65.0% on StrucText-Eval-Hard. In contrast, human participants reached 95.7% accuracy on StrucText-Eval-Hard, underscoring the inadequacy of existing LLMs in processing complex structural information. Utilizing StrucText-Eval data enhances the capabilities of current closed-source LLMs and improves open-source LLMs by fine-tuning them with this benchmark. This finding confirms StrucText-Eval’s critical role in evaluating and optimizing model performance. However, more than using the training data from StrucText-Eval alone is required to surpass GPT4-Turbo’s performance on this test suite. Additionally, not all tasks can be improved solely by training on the relevant dataset. Therefore, StrucText-Eval provides a comprehensive evaluation framework for assessing LLMs’ ability to understand structured text, leaving room for further improvement.
We present the following contributions to our research: (1) We developed a method for generating evaluation data for eight structured languages. This method allows control over the complexity of the generated evaluation data by manipulating the quantity of structural nesting. (2) We introduce StrucText-Eval, a comprehensive benchmark comprising 6,032 samples across eight structured languages and 29 evaluation tasks. This benchmark is designed to assess LLMs’ capabilities in understanding structured data comprehensively. (3) We have conducted extensive experiments, including evaluations of existing large models’ capabilities in understanding structure-rich text and enhancements to the capacity of open-source LLMs in this area. Our findings indicate a significant gap between current LLMs and humans in understanding structure-rich text.
@@ 2.2 Structural Text Understanding Evaluation Datasets
The similar datasets dedicating to structure information understanding are scarce. GraphQA [9] is a benchmark for assessing LLMs on reasoning over graph data, composed of 7 simple graph tasks, with exploration on methods in encoding graph-structured data as text. Struc-Bench [25] collect data from open source to evaluate if LLM can generate complex structured tabular text. Sui et al. [21] consists of 8 tasks on table-formatted input texts, i.e., input data organized in row-column structures, with studying in input design and prompt enhancement. Pasupat and Liang [17] exposed a dataset of 22,033 complex questions on Wikipedia tables and proposed a parsing algorithm for answering complex questions on semi-structured tables. For tabular data, Shwartz-Ziv and Armon [20] showed that deep models were not a good solution; Gorishniy et al. [12] designed two specialized baseline architecture for tabular data and concluded that universally superior solution does not exist. Wang et al. [28] is another graph-based problem solving benchmark, with graph structure described with natural language. Fatemi et al. [9] and Sui et al. [21] include comparison of encoding methods for structure information, i.e., graphs and tables, respectively, showing that LLMs performance varies on different encodings and the relatively vacancy on LLM’s understanding of structure-rich texts input.
Our dataset diverges from previous work in two significant ways. Firstly, our tasks exclusively concentrate on analyzing structural information and executing dependent inference, unlike prior approaches [16, 32] that employ structural data primarily as an auxiliary tool for classification or semantic comprehension. StrucText-Eval focus on symbolic structure by erasing content semantics from input texts, preserving only structural details, which renders the texts semantically obscure and the tasks semantically agnostic. Secondly, our benchmark extends beyond traditional graph reasoning and table information retrieval to include a broader spectrum of structure-rich text types. It also incorporates comprehensive structure-based inference tasks that are tailored to specific input formats.
3 StrucText-Eval Construction
@@ 3.3 Statistic Information
StrucText-Eval has assembled a dataset comprising 6,032 samples, with each of the 29 specific tasks for eight languages containing 208 samples. These tasks are divided into “simple” and “hard” categories based on data length and complexity, as depicted in Fig. 3(a). The dataset is partitioned into training and testing sets at a 9:2 ratio, which supports extensive training while also enabling the evaluation of model generalization on significant volumes of unseen data. Detailed statistics regarding the number of samples, lengths, and complexity levels across all tasks, languages and difficulties are detailed in Tab. 3 in Sec. C in Appendix.
@@ 4 Experiment Setup
To evaluate LLMs’ current capability of processing structure-rich text and executing dependent inference, we conducted a series of experiments using StrucText-Eval in various settings. Our study utilizes both prompt-based and finetuning methods to analyze the performance variations.
@@ 4.3 Finetune-based Method
We further explored the efficacy of finetuning OpenLLaMA 7B v1 using various subsets of data. Our finetuning configurations are described as follows: Base: Utilizes the core StrucText-Eval dataset, referred to as the base StrucText-Eval setting. 2-Shot: Implements a few-shot (2-shot) learning strategy within the StrucText-Eval framework. Rule Hint: Adapts StrucText-Eval by appending a new task-specific rule hint to each sample, effectively redefining the input text as a combination of the question, answer, and original input, with the answer delineated as a step-by-step procedure. Mix: is a variant of Base with half of samples replaced by samples from ShareGPT111https://huggingface.co/datasets/shibing624/sharegpt_gpt4.
@@ 5.1 Close Sourced LLMs’ Performance
In this section, we analyze the performance of closed-source LLMs in various tasks involving structure-rich text and multiple languages. Table 5 displays the overall performance on the complete StrucText-Eval dataset, while Table 2 compares both LLM and human performance specifically on the StrucText-Eval-Hard subset.
@@ 5.3 Case Study
We illustrate the assessment setup and specific evaluation methods of StrucText-Eval through two exemplary cases in Fig. 8.
@@ 6 Conclusion
We believe that the capability to directly interpret structural-rich text in a free-text format is an essential skill all LLMs require. In response, we have developed StrucText-Eval to evaluate this capability of LLMs. Our findings indicate that the proficiency of current LLMs in training on these structural-rich texts varies depending on user frequency, leading to markedly different outcomes when the same tasks are performed in various languages. Currently, LLMs’ understanding of structural-rich texts remains superficially tied to the training data, and these models lack a profound understanding of the structure itself. This deficiency becomes evident when LLMs encounter complex structures composed of common languages or need to parse structural-rich text by custom languages, resulting in significant performance degradation. Direct training on the dataset provided by StrucText-Eval effectively enhances the LLM’s understanding of structural-rich text. However, training solely on the StrucText-Eval dataset does not allow achieving GPT4-Turbo’s comprehensive capabilities on the test set, necessitating the collection of more extensive data and the optimization of training methodologies.
@@ 7 Limitation
This paper focuses on evaluating and enhancing the comprehension LLM’s abilities on structure-rich text by designing a benchmark called StrucText-Eval. However, StrucText-Eval includes only eight types of structured languages and encompasses a total of 29 different tasks. Given the vast array of actual structured languages and the myriad methodologies employed beyond these 29 types, StrucText-Eval can only partially represent the LLMs’ capacity to understand structure-rich text. Additionally, due to regional restrictions, we are unable to utilize some highly effective baseline LLMs, such as Gemini and Claude. Therefore, the conclusions drawn in this paper are based on the assumption that GPT-4 and GPT-4 Turbo represent the top-tier LLMs now.
@@ 8 Ethical Concern
Nature of StrucText-Eval Content: StrucText-Eval is primarily composed of structured language syntax and some nonsensical strings, which do not present potential ethical issues such as gender bias or racial discrimination.
@@
Objective Presentation of Experimental Results: The experimental results pertaining to StrucText-Eval objectively demonstrate the comprehension abilities of various large models on structure-rich text included in the benchmark. We have thoroughly validated the outputs and assessment details of the models to ensure that the entire evaluation adheres to the experimental setup and maintains objectivity.
@@ Appendix A Detail about Manual Works
This paper involves the manual works in writing Question Templates and the acquisition of human performance on StrucText-Eval-Hard. All manual works are carried out by the authors of this paper, so there is no payment for the works. The three authors collectively completed the writing and verification of 29 Question Templates, and all templates along with the dataset have been made publicly available on the same website. Moreover, each of the three authors provided responses to 500 identical questions in StrucText-Eval-Hard, with each author dedicating approximately 17 hours. Thus, the human performance results presented in Table 2 are calculated based on the average scores across these 1500 responses.
Appendix B Statistics to StrucText-Eval
The detailed statistic information about StrucText-Eval is illustrated in Tab. 3.